OpenStudy (anonymous):

Please help ASAP!! How many combinations of 3 students can a teacher choose from 32 students? A. 96 B. 4960 C. 29,760 D. 32,768

OpenStudy (anonymous):

I would say there are \[\left(\begin{matrix}32 \\ 3\end{matrix}\right)\] combos

OpenStudy (anonymous):

It's not C for sure...Can you walk me through it??

OpenStudy (anonymous):

oh oh It's D

OpenStudy (anonymous):

32*32*32

OpenStudy (anonymous):

so there are 32 students lets call them x1, x2, ..... , x32 and you want to make groups of 3 people out of those 32 people it can be {x1, x2, x3}, or {x1, x3, x20}, or {x1, x3, x23} and much more and you want to count all of those possible combos do you know how to expand the equation what i wrote before?

OpenStudy (anonymous):

Umm sort of... So my solution how i got to the answer is wrong im guessing.

OpenStudy (anonymous):

just solve \[\left(\begin{matrix}32 \\ 3\end{matrix}\right) = \frac{ 32! }{ 3!29! }\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

you know how the factorial works right? \[\frac{ 32! }{ 3!29! } = \frac{ (32)(31)(30)....(1) }{ [(3)(2)(1)] [(29)(28)(27)...(1)]}\]

OpenStudy (anonymous):

I kind of know how it works it just using it that gets me stuck

OpenStudy (anonymous):

if you write out the whole thing you can see you can cancel stuff out and you are left with \[\frac{ (32)(31)(30) }{ (3)(2)(1) }\]

OpenStudy (anonymous):

you can simplify it even more and you are left with \[(32)(31)(5)\] and solve that and its your answer

OpenStudy (anonymous):

B

OpenStudy (anonymous):

correct hope you understood what i did